# Learning to Generate Unit Tests for Automated Debugging

### v2

@@ Acknowledgments
We thank Swarnadeep Saha for insightful discussions in the early phases of this work. This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, and the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.
@@ Appendix A Experimental Setup Details
A.1 Coding Datasets
We use debugging datasets based on popular LLM coding benchmarks, HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) along with their extended evaluation suites with more unit tests as proposed by Liu et al. (2024b).
HE+Fix. We start with the HumanEvalFix dataset containing human-introduced errors to gold solutions proposed by Muennighoff et al. (2024) (released under MIT license). However, this dataset only uses minimal unit tests to evaluate code correctness (from HumanEval) which has shown to be unreliable Li et al. (2023), as it can miss errors due to low coverage. Therefore, we filter for problems that overlap with EvalPlus Liu et al. (2024b) released under Apache-2.0 license, which contains over 80 more unit tests. This yields 158 problems that each have a faulty solution and an expanded suite of private UTs – i.e., UTs not used for debugging – for evaluation.
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark Austin et al. (2021); Liu et al. (2024b); Ni et al. (2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Sec. A.1). Then, we select one faulty solution at random corresponding to each problem. This results in a total of 325 problems with incorrect solutions representing realistic errors made by LLM coders. We find such errors are generally more challenging for models to debug (cf. Sec. 5).
@@ A.1 Coding Datasets
HE+Fix. This dataset contains a total of 158 problems each with one incorrect human-written code and has an initial pass rate (prior to debugging) of 0%, i.e., all private unit tests are failing. As mentioned in Sec. 4, we use the dataset provided by Muennighoff et al. (2024)444https://huggingface.co/datasets/bigcode/humanevalpack but replace the test set for each problem from the original test suite in HumanEval Chen et al. (2021) to that in the EvalPlus evaluation suite Liu et al. (2024b). This increases the average unit tests per problem from 8.17 to 775.18 gold unit tests. Note that we have an automatic UT extraction script for the test code in EvalPlus, and we only retain problems for which this extraction is successful (158 out of 164).
MBPP+Fix and MBPP+Fix (Hard). We begin with 378 problems in the MBPP+ dataset Liu et al. (2024b)555https://huggingface.co/datasets/evalplus/mbppplus and follow the same gold UT extraction step described above, discarding problems for which the extraction fails. This leaves us with 375 problems, for which we sample 16 solutions per problem across multiple LLMs: Gemma-7B-IT Team et al. (2024), Llama3 8B Instruct AI@Meta (2024), Llama3.1 8B Instruct Dubey et al. (2024), Tulu-3 8B SFT Lambert et al. (2024a), DeepSeek 7B coder Guo et al. (2024b), Qwen2.5 Coder 7B Hui et al. (2024). To generate MBPP+Fix, we filter for incorrect solutions (i.e., with at least one failing UT) and then randomly sample one incorrect code per problem. This yields 325 problems in total, each with one faulty code. This dataset has an initial pass rate of 24.21% and an average of 107.45 gold unit tests per problem. In order to construct, MBPP+Fix (Hard), we follow a similar process but select only incorrect solutions which pass 50 - 95% of unit tests. The intuition here is that solutions that are partially correct are often harder to debug than those that are fully incorrect. We then randomly sample one such incorrect solution per problem, yielding a dataset of 170 problems with an initial pass rate of 74.83% and an average of 107.49 gold unit tests.
@@
When using generated UTs for improving code generation, we evaluate on HumanEval+ (HE+) and MBPP+ benchmarks proposed in Liu et al. (2024b) released under Apache-2.0 license. As described above, we filter out problems for which we can extract private UTs, yielding 158 problems in HE+ dataset and 375 problems in MBPP+ dataset. For the RM baseline we use the latest version of the 8B model based on Llama-3.1 backbone in Liu et al. (2024a).666https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B-v0.2 We made this choice based on on rankings on the RewardBench leaderboard Lambert et al. (2024b)
@@ A.2 Evaluation Metrics
Code Accuracy. To evaluate the utility of generated unit tests via code debugging, we follow prior work Chen et al. (2023b); Chae et al. (2024) in reporting pass@1 code accuracy, i.e., the percentage of codes passing all unit tests after 3 rounds of debugging. Note that while we debug with model-generated UTs, we evaluate code accuracy using private human-annotated UTs.
@@ Appendix B UTGen Training
Preprocessing Tulu Data. We use the Tulu-3 SFT mixture dataset released by Lambert et al. (2024a) which contains a total of 939.3K instances.777https://huggingface.co/datasets/allenai/tulu-3-sft-mixture However, it contains a mixture of prompts for instruction-following, math reasoning, and coding. Therefore, we filter for prompts involving Python coding by regex search for keywords “python” and “def ” which suit our task setup described in Sec. 3.1. Furthermore, we filter out instances with more than 2K tokens in the prompt and ensure the prompt is a valid unicode string. This results in a total of 48.3K instances for which we use the “prompt” as the problem description and extract code from the response of the last turn when multi-turn interactions are present provided that the extracted code executes without any errors or issues. Finally, we prompt the LLM to be trained with UTGen to generate 2 corrupted versions of this code to serve as the target code, and for each target code, we make 5 attempts to generate failing unit test inputs and filter out instances that do not have any such UTs. This is followed by the rationalization step (with the same LLM to be trained) described in Sec. 3.2 and in Fig. 2, which results in roughly 30K instances for each 7-8B model trained with UTGen and nearly 70K instances for Qwen2.5 32B.
@@
We train each model for 10 epochs with a batch size of 16 and a learning rate of 5e-6 with a cosine learning rate scheduler. Moreover, we only compute negative log-likelihood loss over the completions. We use LoRA Hu et al. (2021) with a rank of 16, = 32 with a dropout of 0.05. We perform checkpoint selection based on the intrinsic Acc. Attack metric. All training and inference is conducted on Nvidia’s A6000 GPUs taking 20 GPU hours for training and GPU hours for inference (generating one code solution and corresponding UTs with test-time scaling). We train the larger Qwen2.5 32B model on 8 A100 GPUs for roughly 40 GPU hours using QLoRa Dettmers et al. (2023) with 4-bit quantization.
@@ Appendix D Additional Intrinsic Evaluation
Appendix E Prompts
In the following pages, we list all the prompts we use in this work.
@@ Appendix E Prompts
[frametitle=Prompted and UTGen Prompt for UT generation] You are given a Python function {signature} to solve the following task:
@@
Input: {wrong_testcase_input}
Output: {wrong_testcase_output}
Expected: {wrong_testcase_expected}
@@
Now **without** coding up the the function check(string), provide step-by-step reasoning for why function check when given input of "BCDFG" generates or returns False.
@@
- The vowels (in lower case) that should be present in the string are: "a", "e", "i", "o" and "u".
- The given input is "BCDEFG" which contains characters: "b", "c", "d", "e", "f", "g".
- While the input string "BCDEFG" contains only vowel "e" and is missing vowels: "a", "i", "o", "u".

### v3

@@ Acknowledgments
We thank Swarnadeep Saha for insightful discussions in the early phases of this work. Furthermore, we thank the anonymous reviewers and area chairs for their helpful feedback throughout the reviewing process. This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, and the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.
@@ Appendix A Experimental Setup Details
A.1 Debugging Datasets
We use debugging datasets based on popular LLM coding benchmarks, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) along with their extended evaluation suites with more unit tests as proposed by (Liu et al., 2024b).
HE+Fix. We start with the HumanEvalFix dataset containing human-introduced errors to gold solutions proposed by Muennighoff et al. (2024) (released under MIT license). However, this dataset only uses minimal unit tests to evaluate code correctness (from HumanEval) which has shown to be unreliable (Li et al., 2023), as it can miss errors due to low coverage. Therefore, we filter for problems that overlap with EvalPlus (Liu et al., 2024b) released under Apache-2.0 license, which contains over 80 more unit tests. This yields 158 problems that each have a faulty solution and an expanded suite of private UTs – i.e., UTs not used for debugging – for evaluation.
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark (Austin et al., 2021; Liu et al., 2024b; Ni et al., 2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Sec. A.1). Then, we select one faulty solution at random corresponding to each problem. This results in a total of 325 problems with incorrect solutions representing realistic errors made by LLM coders. We find such errors are generally more challenging for models to debug (cf. Sec. 5).
@@ A.1 Debugging Datasets
HE+Fix. This dataset contains a total of 158 problems each with one incorrect human-written code and has an initial pass rate (prior to debugging) of 0%, i.e., all private unit tests are failing. As mentioned in Sec. 4, we use the dataset provided by Muennighoff et al. (2024)333https://huggingface.co/datasets/bigcode/humanevalpack but replace the test set for each problem from the original test suite in HumanEval (Chen et al., 2021) to that in the EvalPlus evaluation suite (Liu et al., 2024b). This increases the average unit tests per problem from 8.17 to 775.18 gold unit tests. Note that we have an automatic UT extraction script for the test code in EvalPlus, and we only retain problems for which this extraction is successful (158 out of 164).
MBPP+Fix and MBPP+Fix (Hard). We begin with 378 problems in the MBPP+ dataset (Liu et al., 2024b)444https://huggingface.co/datasets/evalplus/mbppplus and follow the same gold UT extraction step described above, discarding problems for which the extraction fails. This leaves us with 375 problems, for which we sample 16 solutions per problem across multiple LLMs: Gemma-7B-IT (Team et al., 2024), Llama3 8B Instruct (AI@Meta, 2024), Llama3.1 8B Instruct (Dubey et al., 2024), Tulu-3 8B SFT (Lambert et al., 2024a), DeepSeek 7B coder (Guo et al., 2024b), Qwen2.5 Coder 7B (Hui et al., 2024). To generate MBPP+Fix, we filter for incorrect solutions (i.e., with at least one failing UT) and then randomly sample one incorrect code per problem. This yields 325 problems in total, each with one faulty code. This dataset has an initial pass rate of 24.21% and an average of 107.45 gold unit tests per problem. In order to construct, MBPP+Fix (Hard), we follow a similar process but select only incorrect solutions which pass 50 - 95% of unit tests. The intuition here is that solutions that are partially correct are often harder to debug than those that are fully incorrect. We then randomly sample one such incorrect solution per problem, yielding a dataset of 170 problems with an initial pass rate of 74.83% and an average of 107.49 gold unit tests.
@@
When using generated UTs for improving code generation, we evaluate on HumanEval+ (HE+) and MBPP+ benchmarks proposed in Liu et al. (2024b) released under Apache-2.0 license. As described above, we filter out problems for which we can extract private UTs, yielding 158 problems in HE+ dataset and 375 problems in MBPP+ dataset. For the RM baseline we use the latest version of the 8B model based on Llama-3.1 backbone in Liu et al. (2024a).555https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B-v0.2 We made this choice based on on rankings on the RewardBench leaderboard Lambert et al. (2024b)
@@ A.2 Evaluation Metrics
Code Accuracy. To evaluate the utility of generated unit tests via code debugging, we follow prior work (Chen et al., 2023b; Chae et al., 2024) in reporting pass@1 code accuracy, i.e., the percentage of codes passing all unit tests after 3 rounds of debugging. Note that while we debug with model-generated UTs, we evaluate code accuracy using private human-annotated UTs.
@@ Appendix B UTGen Training
Preprocessing Tulu Data. We use the Tulu-3 SFT mixture dataset released by Lambert et al. (2024a) which contains a total of 939.3K instances.666https://huggingface.co/datasets/allenai/tulu-3-sft-mixture However, it contains a mixture of prompts for instruction-following, math reasoning, and coding. Therefore, we filter for prompts involving Python coding by regex search for keywords “python” and “def ” which suit our task setup described in Sec. 3.1. Furthermore, we filter out instances with more than 2K tokens in the prompt and ensure the prompt is a valid unicode string. This results in a total of 48.3K instances for which we use the “prompt” as the problem description and extract code from the response of the last turn when multi-turn interactions are present provided that the extracted code executes without any errors or issues. Finally, we prompt the LLM to be trained with UTGen to generate 2 corrupted versions of this code to serve as the target code, and for each target code, we make 5 attempts to generate failing unit test inputs and filter out instances that do not have any such UTs. This is followed by the rationalization step (with the same LLM to be trained) described in Sec. 3.2 and in Fig. 2, which results in roughly 30K instances for each 7-8B model trained with UTGen and nearly 70K instances for Qwen2.5 32B.
@@
We train each model for 10 epochs with a batch size of 16 and a learning rate of 5e-6 with a cosine learning rate scheduler. Moreover, we only compute negative log-likelihood loss over the completions. We use LoRA (Hu et al., 2021) with a rank of 16, = 32 with a dropout of 0.05. We perform checkpoint selection based on the intrinsic Acc. Attack metric. All training and inference is conducted on Nvidia’s A6000 GPUs taking 20 GPU hours for training and GPU hours for inference (generating one code solution and corresponding UTs with test-time scaling). We train the larger Qwen2.5 32B model on 8 A100 GPUs for roughly 40 GPU hours using QLoRa (Dettmers et al., 2023) with 4-bit quantization.
@@ Appendix D Additional Intrinsic Evaluation
| Intrinsic Metric | Somers’ Delta |
| Attack Rate | 0.43 |
| Output Acc. | 0.34 |
| Acc. Attack (Both) | 0.63 |
To further validate our intrinsic evaluation metrics, specifically, that both high input attack rate and high output accuracy are key to downstream utility of model-generated unit tests – we measure the correlation between each property of the UT and whether the UT yielded successful debugging. To this end, we collect oracle metadata on all unit tests generated for MBPP+Fix with Qwen2.5 7B, i.e., we record if the unit test caused the target code to fail, whether its output was correct, and compute its Somers’ D correlation with whether the target code was debugged successfully for UTs and 1 round of debugging in Table 7. While individually, failing UTs and output correctness show mild correlation with downstream debugging, Table 7 shows that the presence of both attributes yields a strong (and the highest) correlation with debugging success and reinforces the motivation for UTGen and improving unit-test generation ability of LLMs at large.
Appendix E Additional Analysis and Results with UTDebug
On Test-time Scaling and Backtracking.
| UT Method | Acc. | |
| Randomly-sampled (Qwen2.5 7B) | 34.77 | |
| - Output Test-time Scaling | 30.77 | - 4.0 |
| - Backtracking | 32.61 | - 2.2 |
| UTGen (Qwen2.5 7B) | 37.54 | |
| - Output Test-time Scaling | 26.15 | - 11.4 |
| - Backtracking | 34.38 | - 3.2 |
We show that test-time scaling and backtracking are crucial for LLM debugging with generated UTs. From Table 8, we observe that irrespective of the underlying method for UT generation (either randomly sampled or from UTGen), removing either backtracking or test-time scaling hurts downstream performance. First, we find that removing test-time scaling for predicting the output of UTs decreases the performance of randomly-sampled UTs by 4% and that of UTGen by 11.4%. Note that a larger drop in UTGen’s performance when removing test-time scaling of UT output prediction is consistent with the findings in Sec. 5.1 that models struggle with reasoning over correct outputs for failing UT inputs (more often generated by UTGen). Therefore, test-time scaling in output prediction provides a greater boost for reasoning over challenging inputs (Wang et al., 2022), and consequently, removing it yields larger drops for UTGen. Moreover, Table 8 demonstrates that, without validation on other generated unit tests, LLMs tend to overfit on the unit test contained in the feedback, resulting in up to 3.2% drop in performance of UTGen without backtracking.
Comparison with CodeT (Chen et al., 2023a) for Debugging.
| Model | UT Method | HE+Fix | MBPP+Fix | (Hard) |
| Qwen 2.5 7B | No UT | 52.53 | 23.08 | 16.47 |
| Random | 79.75 | 34.77 | 17.06 | |
| Prompted | 75.32 | 32.92 | 24.12 | |
| CodeT | 81.65 | 34.30 | 21.18 | |
| UTGen | 82.91 | 37.54 | 29.41 | |
| Qwen 2.5 32B | No UT | 79.11 | 39.08 | 32.94 |
| Random | 84.81 | 49.54 | 22.94 | |
| Prompted | 85.44 | 50.77 | 40.59 | |
| CodeT | 86.71 | 46.61 | 29.41 | |
| UTGen | 88.61 | 54.15 | 45.29 | |
As mentioned in Sec. 2, UTGen differs from Chen et al. (2023a) in that CodeT is designed for code generation by selecting a code from the largest consensus set of codes and independently generated UT (without conditioning on code). Alternatively, this UT sampling procedure is akin to the randomly-sampled baseline (described in Sec. 4) but the latter uses self-consistency for predicting the UT output. In contrast, UTGen is designed for UT generation focusing on the intrinsic quality of UTs as well as their utility for downstream tasks like code-debugging and best-of-N ranking by conditioning UT generation on the edge cases of a target code. To directly compare against CodeT, instead of using the code generated from CodeT, we use its unit tests in downstream debugging. In Table 9, we scale up CodeT such that it uses a similar computational budget as UTGen and sample unit tests for debugging with Qwen 2.5 7B and 32B Code-Instruct models. Across both models and all three debugging datasets, UTGen outperforms debugging with unit tests generated by CodeT by as much as 7.54% (absolute) on MBPP+Fix and 15.88% (absolute) on MBPP+Fix (Hard) with Qwen 2.5 32B model. Moreover, CodeT (which generates UTs independent of code being debugged) lags behind the prompted UT baseline in nearly half of the settings, showing that conditioning UT generation on the erroneous code better helps identify and localize bugs.
Appendix F Scaling with Number of Generated UTs
Appendix G Discussion
Our work on UTGen contributes to the broader landscape of verification and feedback generation for LLM-generated code. While recent work has focused on training verifiers to provide feedback (Mahan et al., 2024; Zhang et al., 2024), a key challenge remains in obtaining high-quality feedback signals for debugging. UTGen addresses this by directly generating unit tests that can identify problems in code, complementing existing work on how to effectively incorporate and present feedback for debugging (Chen et al., 2023b; Zhong et al., 2024) along with test-time scaling and backtracking incorporated in UTDebug. Our results demonstrate that without a quality signal to determine code correctness and/or how a faulty code is failing (in the form of unit tests), using LLMs to generate feedback and debug still proves to be challenging. This is one of the first efforts in this direction, and we hope to spark more interest in future work toward LLM-generated unit tests (both input and outputs) that reveal the model’s coding errors.
Our approach connects to and complements recent work on handling real-world programming issues. While approaches designed for SWEBench (Jimenez et al., 2024) focus on fixing known issues from GitHub by understanding and implementing fixes for bug reports, UTGen addresses the upstream challenge of automatically discovering potential issues in new code through test generation. Both tasks share a core challenge: determining the expected behavior of code without access to correct implementations. This connects to the fundamental concept of simulatability from computability theory (Rice, 1953), where we ask whether a program can predict the behavior of another program. Recent work such as Jain et al. (2024) shows that while LLMs can often simulate existing code by tracing execution steps, they struggle more with predicting correct outputs from specifications alone. Our results align with these findings – while UTGen can generate test inputs that trigger errors (high attack rate), predicting correct expected outputs remains challenging (lower output accuracy). This suggests that improving LLMs’ ability to reason about intended program behavior from specifications remains a crucial direction for future work. Nevertheless, we find that the modifications made to debugging in UTDebug help boost UTGen’s accuracy and account for noise, leading to downstream gains.
Appendix H Prompts
@@
In the following pages, we list all the prompts we use in this work. {user_example}[frametitle=Prompted and UTGen Prompt for UT generation] You are given a Python function {signature} to solve the following task:
@@
Input: {wrong_testcase_input}
Output: {wrong_testcase_output}
Expected: {wrong_testcase_expected}
@@
Now **without** coding up the the function check(string), provide step-by-step reasoning for why function check when given input of ”BCDFG” generates or returns False.
@@
- The vowels (in lower case) that should be present in the string are: ”a”, ”e”, ”i”, ”o” and ”u”.
- The given input is ”BCDEFG” which contains characters: ”b”, ”c”, ”d”, ”e”, ”f”, ”g”.
- While the input string ”BCDEFG” contains only vowel ”e” and is missing vowels: ”a”, ”i”, ”o”, ”u”.